# Comparison of renormalized interactions using one-dimensional few-body systems as a testbed

### v1

@@ IV.2 Density
Figure 4: Panels a) and b): Density of the ground state calculated with the potentials $g\delta(x)$ and $g(n)\delta(x)$ and two cutoff parameters. Panels c) and d): Density of the ground state in the center of the trap as a function of the cutoff parameter, $n$. Panels a) and c) show results for the $1+1$ system; panels b) and d) are for $1+3$. All panels are for attractive interactions: $g=-2.5067$, which corresponds to $E_{1+1}=-1$. Numerical data in panels c) and d) are shown with markers of different shapes. Lines are added to guide the eye. In the supplementary material [56], we also present these data as a function of $1/n$ and using the log-log scale for $1+1$.
In Fig. 4 we show the ground-state density for attractive interactions for the $1+1$ [panels a) and c)] and $1+3$ systems [panels b) and d)]. The two upper panels show the density calculated with the bare contact interaction and with the running coupling constant approach. The two lower panels show the density at the trap center for all interactions. The plot demonstrates that the density converges fast to the limit $n\to\infty$ for $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ even without transforming the density operator alongside the interaction as discussed Sec. III.2. Therefore, we leave the study of the effect of this transformation on the density to future works [^10] Note that the induced two-body terms (see Eq. (17)) $\rho^{(2)}(x_{\alpha},x_{\beta})=\sum_{\alpha,\beta=1}^{N}\left(\rho(x_{\alpha})+\rho(x_{\beta})\right)$ (26) do not factorize in relative $x_{\alpha,\beta}=\frac{1}{\sqrt{2}}(x_{\alpha}-x_{\beta})$ and center-of-mass $X_{\alpha,\beta}=\frac{1}{\sqrt{2}}(x_{\alpha}+x_{\beta})$ coordinates, which complicates their computation. For more details see the supplementary material [56]..
While the overall shape of the density is reproduced with all interactions and even for a small cutoff, large values of $n$ are still required for an accurate estimate of the density. We illustrate this by considering the density at the center of the trap, $\rho(x=0)$, see panels c) and d). For the $1+1$ system we see that results obtained with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ approach the exact value quickly. Similar to the energy, the potentials $g(n)\delta(x)$ and $V_{\mathrm{eff}}^{\mathrm{rel}}$ enjoy similar convergence patterns, and the bare contact potential produces least accurate results for the density. The same conclusion holds also for larger systems, see Fig. 4 d).
@@
The analysis allows us to quantify a quick convergence of the density calculated with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ ($\sigma>1$). For a detailed description of the fitting analysis see App. B. This shows that although we optimized the interactions for the energies, they also lead to better results for the density, in agreement with previous studies [20, 34, 37].
@@ IV.3 Transformation of observables
In this section, we focus on the transformation of operators together with the potential (see Eq. (17)), which has received little attention in the corresponding cold-atom literature. To this end, we calculate the expectation values of the one-body parts of the Hamiltonian, i.e., of the kinetic energy operator [^11] Note that strictly speaking the kinetic energy is not an observable quantity. In one spatial dimension it is well-defined, however, in higher dimensions its value diverges. This can be shown, e.g., using the Tan’s contact parameter [62]. Nevertheless, we discuss the convergence of $\langle T\rangle$ as $T$ is one of the simplest operators that is sensitive to changes of the wave function on small length scales., $T$, and of the trapping potential, $W$, see Eq. (2). The effective interactions $V_{\mathrm{eff}}^{\mathrm{rel}}$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ are constructed via a unitary transformation from which the transformation of operators alongside the construction of these potentials follows. The running coupling constant approach cannot be represented via a unitary transformation. Therefore, we cannot transform the operator together with this potential, and only present the expectation values of non-transformed operators in this case. Similar to the energy, the relevant scale to judge the convergence is the harmonic oscillator energy unit, $\hbar\omega=1$.
@@
Figure 5: Expectation value of the kinetic energy operator, $\langle T\rangle$, as a function of the cutoff parameter, $n$. Left panels [*1)] show $\langle T\rangle$ for the operator $T$ that has been transformed together with $V_{\mathrm{eff}}^{\mathrm{rel}}$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. Right panels [*2)] show what happens if the potential is renormalized but the operator $T$ is not transformed. Panels a) and b) are for the $1+1$ system, while panels c) and d) are for $1+3$. Panels a) and c) are for $g=-2.5067$ (corresponds to $E_{1+1}=-1$); panels b) and d) are for $g=3$ (corresponds to $E_{1+1}=1.6$). Our data are shown with markers; lines are added to guide the eye. We show these data in the supplementary material [56] also as a function of $1/n$ and (for the $1+1$ system) using the log-log scale.
@@
Figure 6: Expectation value of the trap operator, $\langle W\rangle$, as a function of the cutoff parameter, $n$. Left panels [*1)] show $\langle W\rangle$ for the operator $W$ that has been transformed together with $V_{\mathrm{eff}}^{\mathrm{rel}}$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. Right panels [*2)] show what happens if the potential is renormalized but the operator $W$ is not transformed. Panels a) and b) are for the $1+1$ system, while panels c) and d) are for $1+3$. Panels a) and c) are for $g=-2.5067$ (corresponds to $E_{1+1}=-1$); panels b) and d) are for $g=3$ (corresponds to $E_{1+1}=1.6$). Our data are shown with markers; lines are added to guide the eye. In the supplementary material [56], we show these data also as a function of $1/n$ and (for the $1+1$ system) using the log-log scale.
@@ IV.4 Transition matrix elements
assuming that $|\eta|\;\ll|g(0)|$. The transition from the ground state |$\ket{0}$ to an excited state |$\ket{f}$ driven by this modulation can be estimated using Fermi’s golden rule as $R_{0\to f}\propto\eta^{2}|M_{0\to f}|^{2}\delta(E_{0}-E_{f}-\omega)$. The focus of this subsection is on the transition matrix element,
@@
Figure 7: Transition matrix elements from the ground state to the lowest-energy state that can be excited by a periodic modulation of the interaction. Left panels [*1)] show $M_{0\to f}$ for the observable that has been transformed together with the potential while the right ones [*2)] show what happens if the potential has been renormalized but the operator has not been modified. Panels a) and b) are for the $1+1$ system while panels c) and d) are for the $1+3$ system. Panels a) and c) are for $g=-2.5067$ (corresponds to $E_{1+1}=-1$); panels b) and d) are for $g=3$ (corresponds to $E_{1+1}=1.6$). Our data are shown with markers; lines are added to guide the eye. In the supplementary material [56], we show these data also as a function of $1/n$ and (for the $1+1$ system) using the log-log scale.
@@
The discussion in this subsection highlights both the necessity of renormalizing interactions and the importance of transforming operators, such as the transition matrix elements, in numerical studies of few-fermion models. In particular, we are able to obtain accurate numerical data for attractive $1+3$ systems with only $n\simeq 10$, which leads to matrices that can be diagonalized on a standard PC.
@@ V.1 Summary
We have contrasted zero-range one-dimensional interactions useful in studies of a few cold atoms in a trap. In agreement with previous studies, we have demonstrated that numerically exact solutions to a few-body problem can be obtained by explicitly including information about converged two-body energies into the calculations in a truncated Hilbert space.
A comparison between the renormalization schemes has demonstrated that a simple running coupling constant approach (used in particular in Refs. [34, 37]) might be a go-to method for approximate calculations of the energy spectrum as it leads to a convergence pattern similar to $V_{\mathrm{eff}}^{\mathrm{rel}}$ also used in previous studies [20, 32, 30]. Indeed, running coupling constant approach is simple and transparent as it relies on a one-parameter renormalization of the interaction strength. At the same time, more accurate results at (almost) the same numerical cost can be obtained with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. This effective potential fixes the two-body spectrum in a truncated Hilbert space to that in an infinite Hilbert space, which should be calculated either numerically or analytically. Although this potential has received little attention in the cold-atom literature, we hope that our results will motivate its use in the future.
The advantage of numerical methods based upon $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ and $V_{\mathrm{eff}}^{\mathrm{rel}}$ becomes clear in calculations of observables, as these potentials are generated by a unitary transformation. We have observed that transformation of observables (in particular of the kinetic energy and transition matrix elements) can improve convergence considerably. In particular, we have noticed that our data calculated with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ can be considered numerically exact already for relatively small sizes of the Hilbert space. This is because the harmonic oscillator breaks translational invariance. In the oscillator relative and center-of-mass motion only decouple for infinite basis size. The effective interaction $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ corrects for this finite basis artefact and converges more quickly. For large basis size the relative and center-of-mass motions are effectively decoupled and $V_{\mathrm{eff}}^{\mathrm{rel}}$ gives similar results as $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$.
@@ V.2 Outlook
Our study motivates a comparative study of regularization methods employed in two-dimensional geometries [21, 23, 22], which are explored less than in one and three spatial dimensions. In Fig. 8, we showcase calculations for a two-dimensional harmonic oscillator confiment (see App. C for the Hamiltonian and basis functions) that illustrate that already a simple potential $V_{\mathrm{eff}}^{\mathrm{rel}}$ can boost convergence of the energy. Here, we regularize the diverging two-dimensional $\delta$-potential by fixing the two-body ground state energy for each cutoff [21]. The cutoff is given by the highest one-body energy ($n_{2D}$) of the one-body basis used to construct the Hamiltonian. We show results for the first excited state of $1+1$ with the total angular momentum of zero and the ground-state energy for 1+2121+21 + 2141414We have observed similar convergence patterns also for other particle numbers such as $3+3$ fermions.. We expect that the use of the effective interaction will be beneficial also in calculations of observables, assuming that they are transformed together with the potential.
We emphasize that our discussion on renormalized interaction is not limited to CI calculations and zero-range potentials, and should be useful also in other numerical methods and for non-contact interactions, e.g., between cold dipoles [63, 64, 65]. In an upcoming study, we will combine renormalized interactions and importance truncated CI (see e.g. Refs [45, 25, 24]) to calculate systems with larger particle numbers (mainly enabled by the importance truncation) and strong interactions (mainly enabled by renormalized interactions).
@@ Acknowledgements
We thank J. Cremon and J. Bjerlin for earlier contributions to the configuration interaction calculations used in this work (see Refs. [49, 50]). This work was supported by a fellowship of the German Academic Exchange Service (DAAD), the Swedish Research Council and the Knut and Alice Wallenberg foundation. F.B. and S.M.R. acknowledges helpful discussions with Carl Heintze, Sandra Brandstetter and Lila Chergui. We further want to thank Lila Chergui for helpful comments on the manuscript.
@@ Appendix B Analysis of convergence
The fitting function in Eq. (37) is partly inspired by the discussion in App. A, Eq. (36), where we show the convergence behavior of the energy of many fermions in a one-dimensional harmonic trap interacting via bare contact interaction (see also Ref. [13]). We therefore expect for a bare contact interaction $\sigma\approx 0.5$ for the convergence of the energy. In Ref. [20] a similar convergence rate was postulated and found for the density. In that reference, the authors studied the convergence of the energies and the density at $x=0$ using the effective interaction $V_{\mathrm{eff}}^{\mathrm{rel}}$ and the delta function potential. We extend their study by also considering the running coupling constant and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. We further discuss the convergence rates of the transition matrix elements for a periodic modulation of the interaction strength, the expectation value of the kinetic energy as well as of the trapping potential.
Note, that we do not know the convergence pattern for the data computed using the renormalized interactions. In fact, these data do not have to follow Eq. (37). Therefore, in all fitted data there is a systematic error which we cannot estimate. Nevertheless, fitting the data with Eq. (37) can help us to quantify the convergence, e.g., larger values of $\sigma$ showcase faster convergence.

### v2

@@ IV.2 Density
Figure 4: Panels a) and b): Density of the ground state calculated with the potentials $g\delta(x)$ and $g(n)\delta(x)$ and two cutoff parameters. Panels c) and d): Density of the ground state in the center of the trap as a function of the cutoff parameter, $n$. Panels a) and c) show results for the $1+1$ system; panels b) and d) are for $1+3$. All panels are for attractive interactions: $g=-2.5067$, which corresponds to $E_{1+1}=-1$. Numerical data in panels c) and d) are shown with markers of different shapes. Lines are added to guide the eye. In the supplementary material [56], we also present these data as a function of $1/n$, Fig. S2, and using the log-log scale for $1+1$, Fig. S7.
In Fig. 4 we show the ground-state density for attractive interactions for the $1+1$ [panels a) and c)] and $1+3$ systems [panels b) and d)], whose operator is defined in first quantization picture as
$\rho(z)=\sum\limits_{i=1}^{N}\delta(z-x_{i})$ (26)
while in second quantization it reads
$\rho(z)=\sum\limits_{i,j}\Phi_{i}(z)\Phi_{j}(z)a_{i}^{\dagger}a_{j}\,.$ (27)
The two upper panels show the density calculated with the bare contact interaction and with the running coupling constant approach. The two lower panels show the density at the trap center for all interactions. The plot demonstrates that the density converges fast to the limit $n\to\infty$ for $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ even without transforming the density operator alongside the interaction as discussed Sec. III.2. Therefore, we leave the study of the effect of this transformation on the density to future works [^11] Note that the induced two-body terms (see Eq. (17)) $\rho^{(2)}(z_{\alpha},z_{\beta})=\sum_{\alpha,\beta=1}^{N}\left(\rho(z_{\alpha})+\rho(z_{\beta})\right)$ (28) do not factorize in relative $z_{\alpha,\beta}=\frac{1}{\sqrt{2}}(z_{\alpha}-z_{\beta})$ and center-of-mass $Z_{\alpha,\beta}=\frac{1}{\sqrt{2}}(z_{\alpha}+z_{\beta})$ coordinates, which complicates their computation. For more details see Sec. III in the supplementary material [56]..
While the overall shape of the density is reproduced with all interactions and even for a small cutoff, large values of $n$ are still required for an accurate estimate of the density. We illustrate this by considering the density at the center of the trap, $\rho(z=0)$, see panels c) and d). We have checked that the convergence of the density at the center is representative also of other values of $z$. To this end, we investigated the behavior of the relative density (not shown), defined as $(\rho_{n=11}(z)-\rho_{n=31}(z))/\rho_{n=31}(z)$.
For the $1+1$ system we see that results obtained with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ approach the exact value quickly. Similar to the energy, the potentials $g(n)\delta(x)$ and $V_{\mathrm{eff}}^{\mathrm{rel}}$ enjoy similar convergence patterns, and the bare contact potential produces least accurate results for the density. The same conclusion holds also for larger systems, see Fig. 4 d).
@@
The analysis allows us to quantify a quick convergence of the density calculated with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ ($\sigma>1$). For a detailed description of the fitting analysis see App. B. This shows that although the effective interactions are optimized using the energy spectrum, they also lead to better results for the density, in agreement with previous studies [20, 34, 37].
@@ IV.3 Transformation of observables
In this subsection, we focus on the transformation of operators together with the potential (see Eq. (17)), which has received little attention in the corresponding cold-atom literature. To this end, we calculate the expectation values of the one-body parts of the Hamiltonian, i.e., of the kinetic energy operator [^12] Note that strictly speaking the kinetic energy is not an observable quantity. In one spatial dimension it is well-defined, however, in higher dimensions its value diverges. This can be shown, e.g., using the Tan’s contact parameter [62]. Nevertheless, we discuss the convergence of $\langle T\rangle$ as $T$ is one of the simplest operators that is sensitive to changes of the wave function on small length scales., $T$, and of the trapping potential, $W$, see Eq. (2). The effective interactions $V_{\mathrm{eff}}^{\mathrm{rel}}$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ are constructed via a unitary transformation from which the transformation of operators alongside the construction of these potentials follows. The running coupling constant approach cannot be represented via a unitary transformation. Therefore, we cannot transform the operator together with this potential, and only present the expectation values of non-transformed operators in this case. Similar to the energy, the relevant scale to judge the convergence is the harmonic oscillator energy unit, $\hbar\omega=1$.
@@
Figure 5: Expectation value of the kinetic energy operator, $\langle T\rangle$, as a function of the cutoff parameter, $n$. Left panels [*1)] show $\langle T\rangle$ for the operator $T$ that has been transformed together with $V_{\mathrm{eff}}^{\mathrm{rel}}$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. Right panels [*2)] show what happens if the potential is renormalized but the operator $T$ is not transformed. Panels a) and b) are for the $1+1$ system, while panels c) and d) are for $1+3$. Panels a) and c) are for $g=-2.5067$ (corresponds to $E_{1+1}=-1$); panels b) and d) are for $g=3$ (corresponds to $E_{1+1}=1.6$). Our data are shown with markers; lines are added to guide the eye. We show these data in the supplementary material [56] also as a function of $1/n$, Fig. S3, and (for the $1+1$ system) using the log-log scale, Fig. S8.
@@
Figure 6: Expectation value of the trap operator, $\langle W\rangle$, as a function of the cutoff parameter, $n$. Left panels [*1)] show $\langle W\rangle$ for the operator $W$ that has been transformed together with $V_{\mathrm{eff}}^{\mathrm{rel}}$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. Right panels [*2)] show what happens if the potential is renormalized but the operator $W$ is not transformed. Panels a) and b) are for the $1+1$ system, while panels c) and d) are for $1+3$. Panels a) and c) are for $g=-2.5067$ (corresponds to $E_{1+1}=-1$); panels b) and d) are for $g=3$ (corresponds to $E_{1+1}=1.6$). Our data are shown with markers; lines are added to guide the eye. In the supplementary material [56], we show these data also as a function of $1/n$, Fig. S4 and (for the $1+1$ system) using the log-log scale, Fig. S9.
@@ IV.4 Transition matrix elements
assuming that $|\eta|\;\ll|g(0)|$. The transition from the ground state |$\ket{0}$ to an excited state |$\ket{f}$ driven by this modulation can be estimated using Fermi’s golden rule as $R_{0\to f}\propto\eta^{2}|M_{0\to f}|^{2}\delta(E_{f}-E_{0}-\omega)$. The focus of this subsection is on the transition matrix element,
@@
Figure 7: Transition matrix elements from the ground state to the lowest-energy state that can be excited by a periodic modulation of the interaction. Left panels [*1)] show $M_{0\to f}$ for the observable that has been transformed together with the potential while the right ones [*2)] show what happens if the potential has been renormalized but the operator has not been modified. Panels a) and b) are for the $1+1$ system while panels c) and d) are for the $1+3$ system. Panels a) and c) are for $g=-2.5067$ (corresponds to $E_{1+1}=-1$); panels b) and d) are for $g=3$ (corresponds to $E_{1+1}=1.6$). Our data are shown with markers; lines are added to guide the eye. In the supplementary material [56], we show these data also as a function of $1/n$, Fig. S5, and (for the $1+1$ system) using the log-log scale, Fig. S10.
@@
The discussion in this subsection highlights both the necessity of renormalizing interactions and the importance of transforming operators, such as the transition matrix elements, in numerical studies of few-fermion models. In particular, we are able to obtain accurate numerical data for attractive $1+3$ systems with only $n\simeq 10$, which leads to matrices that can be diagonalized on a standard computer.
@@ V.1 Summary
We have contrasted zero-range one-dimensional interactions useful in studies of a few cold atoms in a trap. In agreement with previous studies, we have demonstrated that numerically exact solutions to a few-body problem can be obtained by explicitly including information about converged two-body energies into the many-body calculations in a truncated Hilbert space.
A comparison between the renormalization schemes has demonstrated that a simple running coupling constant approach (used in particular in Refs. [34, 37]) might be a go-to method for approximate calculations of the energy spectrum as it leads to a convergence pattern similar to $V_{\mathrm{eff}}^{\mathrm{rel}}$ also used in previous studies [20, 32, 30]. Indeed, running coupling constant approach is simple and transparent as it relies on a one-parameter renormalization of the interaction strength. At the same time, more accurate results at (almost) the same numerical cost can be obtained with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. This effective potential fixes the two-body spectrum in a truncated Hilbert space to that in an infinite Hilbert space. The spectrum can be calculated either numerically or analytically. Although this potential has received little attention in the cold-atom literature, we hope that our results will motivate its use in the future.
The advantage of numerical methods based upon $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ and $V_{\mathrm{eff}}^{\mathrm{rel}}$ becomes clear in calculations of observables, as these potentials are generated by a unitary transformation. We have observed that transformation of observables (in particular of the kinetic energy and transition matrix elements) can improve convergence considerably. In particular, we have noticed that our data calculated with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ can be considered numerically exact already for relatively small sizes of the Hilbert space. One reason for that follows from the observation that the harmonic oscillator breaks translational invariance. In the oscillator relative and center-of-mass motion only decouple for infinite basis size. The effective interaction $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ corrects for this finite basis artefact and converges more quickly. For large basis sizes, the relative and center-of-mass motions are effectively decoupled and $V_{\mathrm{eff}}^{\mathrm{rel}}$ gives similar results as $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$.
@@ V.2 Outlook
Our study motivates a comparative study of regularization methods employed in two-dimensional geometries [21, 23, 22], which are explored less than in one and three spatial dimensions. In Fig. 8, we showcase calculations for a two-dimensional harmonic oscillator confiment (see App. C for the Hamiltonian and basis functions) that illustrate that already a simple potential $V_{\mathrm{eff}}^{\mathrm{rel}}$ can boost convergence of the energy. Here, we regularize the diverging two-dimensional $\delta$-potential by fixing the two-body ground state energy for each cutoff [21]. The cutoff is given by the highest one-body energy ($n_{2D}$) of the one-body basis used to construct the Hamiltonian. We show results for the first excited state of $1+1$ whose total angular momentum is zero and the ground-state energy for 1+2121+21 + 2151515We have observed similar convergence patterns also for other particle numbers such as $3+3$ fermions.. We expect that the use of the effective interaction will be beneficial also in calculations of observables, assuming that they are transformed together with the potential.
We emphasize that our discussion on renormalized interaction is not limited to CI calculations and zero-range potentials. It can be extended to other numerical methods and to beyond ‘simple’ delta-function potentials, e.g., dipole-dipole [63, 64, 65] or odd-channel interactions [66, 67, 5]. In an upcoming study, we will combine renormalized interactions and importance truncated CI (see e.g. Refs [45, 25, 24]) to calculate systems with larger particle numbers (mainly enabled by the importance truncation) and strong interactions (mainly enabled by renormalized interactions).
@@ Acknowledgements
We thank J. Cremon and J. Bjerlin for earlier contributions to the configuration interaction calculations used in this work (see Refs. [49, 50]). F.B. and S.M.R. acknowledges helpful discussions with Carl Heintze, Sandra Brandstetter and Lila Chergui. We further want to thank Lila Chergui for helpful comments on the manuscript. This work was supported by a fellowship of the German Academic Exchange Service (DAAD), the Swedish Research Council and the Knut and Alice Wallenberg foundation.
@@ Appendix B Analysis of convergence
The fitting function in Eq. (39) is partly inspired by the discussion in App. A, Eq. (38), where we show the convergence behavior of the energy of many fermions in a one-dimensional harmonic trap interacting via bare contact interaction (see also Ref. [13]). We therefore expect for a bare contact interaction $\sigma\approx 0.5$ for the convergence of the energy. In Ref. [20] a similar convergence rate was postulated and found for the density. In that reference, the authors studied the convergence of the energies and the density at $z=0$ using the effective interaction $V_{\mathrm{eff}}^{\mathrm{rel}}$ and the delta function potential. We extend their study by also considering the running coupling constant and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. We further discuss the convergence rates of the transition matrix elements for a periodic modulation of the interaction strength, the expectation value of the kinetic energy as well as of the trapping potential.
Note, that we do not know the convergence pattern for the data computed using the renormalized interactions. In fact, these data do not have to follow Eq. (39). Therefore, in all fitted data there is a systematic error which we cannot estimate. Nevertheless, testing the validity of our fits with a $\chi^{2}$-test showed a chi-squared test statistic smaller than $10^{-8}$ and a p-value of one, which shows that Eq. (39) models our data adequately. We conclude that fitting the data with Eq. (39) helps one to quantify the convergence, e.g., larger values of $\sigma$ showcase faster convergence.